A web-based store is stuck with a very large set of three kinds of unsold Christmas snow globes: Globes with a Nutcracker figurine, globes with a Santa figurine, and globes with a Snowman (see pictures below from left to right). 2.o Fifty percent of the snow globes have the Nutcracker, 20% have Santa, and 30% have the Snowman. The store runs a special where they offer 10 randomly chosen snow globes for $100-a roughly 10 for the price of 4 special. (3 pts) Define random variables for each of the types of snow globes and give the name and parameters of the joint distribution of these random variables. a) b) (3 pts) Find the probability of getting 3 Nutcracker snow globes and 7 Snowman snow globes in a set of 10 snow globes. Show all work. (6 pts) Let X= three times the number of nutcrackers plus two times the number of snowmen. the mean and variance of X c) Find
Answer #1

a) Out of n=10 randomly chosen snow globes, let

  • egin{align*}X_1end{align*} indicate the number of Nutcracker snow globes
  • egin{align*}X_2end{align*} indicate the number of Santa snow globes
  • egin{align*}X_3end{align*} indicate the number of Snowman snow globes

The joint distribution of egin{align*}X_1,X_2,X_3end{align*} has a multinomial distribution with parameters n=10 trials and probabilities p1 = 0.50. P2 = 0.20,P3 = 0.30

the joint pmf is given by

n! 7l 10! 0.5011 × 0.2012 × 0.3013

Formally the joint pmf  of egin{align*}X_1,X_2,X_3end{align*} is

10! -0.50ェ1x 0.20°2 0.3013 , xi,T2,T3-0, 1.2, , 10: x1+x2 +2.3-10 r12

b) The probability of getting 3 Nutcracker snow globes, 0 Santa snow globes and 7 Snowman snow globes in a set of 10 snow globes is

P(3,0,7) 101-0503 ×0.20#x0.30. = 3!0!7! 10 × 9 × × 1 (3x2x1)×(1) × (7 x6x.M)0.503 x 1 x 0.30 (3 × 2 × 1) × (1) × (7 × 6 × × 1) 10×9×8 0.503 × 1 × 0.307 -(3 × 2 × 1) 0.0032805

The probability of getting 3 Nutcracker snow globes, and 7 Snowman snow globes in a set of 10 snow globes is 0.0033

c) Let egin{align*}X_1end{align*} indicate the number of Nutcracker snow globes out of 10 randomly selected globes. The marginal pmf of egin{align*}X_1end{align*} is a Binomial distribution with parameters number of trails, n=10 and probability of success p=0.50

The expected value of egin{align*}X_1end{align*} is (using the result for Binomial distribution)

E( Xi) = n × p= 10 x 0.50-5

The variance of egin{align*}X_1end{align*} is

Var(X) = n × p × (1-p) 10 × 0.50 × (1-0.50) = 2.5

Let egin{align*}X_3end{align*} indicate the number of Snowman snow globes out of 10 randomly selected globes. The marginal pmf of egin{align*}X_3end{align*} is a Binomial distribution with parameters number of trails, n=10 and probability of success p=0.30

The expected value of egin{align*}X_3end{align*} is (using the result for Binomial distribution)

E(X) = n × p= 10 x 0.30 = 3

The variance of egin{align*}X_3end{align*} is

egin{align*} Var(X_3)=n imes p imes (1-p)=10 imes 0.30 imes (1-0.30)=2.1 end{align*}

The covariance of egin{align*}X_1,X_3end{align*} is (using the standard result of a multinomial distribution)

egin{align*} Cov(X_1,X_3)=-np_1p_3=-10 imes 0.5 imes 0.3=-1.5end{align*}

Now egin{align*} X=3X_1+2X_3end{align*}

For any 2 random variables Y, Z and constants a,b we have the following results

egin{align*} E(aY+bZ)=aE(Y)+bE(Z)end{align*}

egin{align*} Var(aY+bZ)=a^2Var(Y)+b^2Var(Z)+2abCov(Y,Z)end{align*}

The expected value of X is

egin{align*} E(X)&=E(3X_1+2X_3)=3E(X_1)+2E(X_3)=3 imes 5+2 imes 3 = 21end{align*}

the variance of X is

egin{align*} Var(X)&=Var(3X_1+2X_3) &=3^2Var(X_1)+2^2Var(X_3)+2 imes 3 imes 2 imes Cov(X_1,X_3) &=3^2 imes 2.5+2^2 imes 2.1+2 imes 3 imes 2 imes (-1.5) &=12.9 end{align*}